OpenStudy (anonymous):

what is the x-intercept of this line y-3=(x+1) I think the answer is x=(-1)

OpenStudy (solomonzelman):

\[y-3=(x+1)\]\[y-3=x+1\]to find the x-intercept set y=0 \[\color{red} {(0) }-3=x+1\]\[-3=x+1\]\[-3\color{red} {-1}=x+1\color{red} {-1}\]\[\color{blue} {-4=x}\]

OpenStudy (solomonzelman):

\[\color{blue} {(-4,~0)}\]is the coordinate.

OpenStudy (anonymous):

thank you so much you explained this perfectly your a life saver I was waaaay off once again thanks a million

OpenStudy (solomonzelman):

Mentor, and close to champion, but I don't really care this much.